OpenStudy (anonymous):

x^(7/6) = 512 what is x?

OpenStudy (anonymous):

x=128 x 2^5/7 approximately 210.01

zepdrix (zepdrix):

\[\Large\rm x^{7/6}=512\] \[\Large\rm \left(x^{7/6}\right)^{6/7}=512^{6/7}\] \[\Large\rm x=512^{6/7}\]

OpenStudy (anonymous):

thanks